Circular No. 8910

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2007 S5–S10 AND C/2007 T7–T11 (SOHO) Further to IAUC 8896, additional Kreutz-sungrazing comets have been

found on SOHO website images, as tabulated below. C/2007 S5 was bright (mag 4) and teardrop-shaped in C3-coronagraph images; C2 images showed a long (0^{o}.75), thin tail. C/2007 S6, C/2007 S8, C/2007 S9, and C/2007 T8 were described by K. Battams as tiny and stellar in appearance (mag ∼ 7, except mag 7.5 for C/2007 S8) in C3 images; C2 images showed C/2007 S6 and C/2007 S8 to be very faint and diffuse, while C/2007 S9 was “fuzzy” and faint, and C/2007 T8 was diffuse. C/2007 S7 appeared very diffuse and faint (mag 8). C/2007 S10 and C/2007 T7 were also small and stellar in appearance (mag ∼ 6) in C3 images, with the former being diffuse and the latter a bit diffuse with a hint of tail in C2 images. C/2007 T9 (also found by R. Matson and T. Hoffman) was small and stellar in appearance (mag 6.5) in C3 images, and diffuse with a hint of tail in C2 images. C/2007 T10 was stellar in appearance (mag 7) in C3 images, and diffuse in C2 images. C/2007 T11 was very faint (mag ∼ 8) and diffuse.

Comet 2007 UT α_{2000} δ_{2000} Inst. F MPEC
C/2007 S5  Sept. 26.654 11^{h}41.^{m}2 − 2^{o}48′ C3/2 RM 2008-A40
C/2007 S6  28.263 12 04.0 − 2 39 C3/2 HS 2008-A40
C/2007 S7 28.704 12 11.4 − 2 44 C2 HS 2008-A40
C/2007 S8  29.263 12 05.4 − 3 11 C3/2 HS 2008-A40
C/2007 S9  29.471 12 06.0 − 3 25 C3/2 HS 2008-A40
C/2007 S10  30.013 12 02.8 − 3 57 C3/2 BZ 2008-A40
C/2007 T7 Oct. 4.696 12 22.7 − 5 44 C3/2 RM 2008-A41
C/2007 T8  10.471 12 45.1 − 7 58 C3/2 RM 2008-A41
C/2007 T9  10.821 12 44.5 − 8 19 C3/2 BZ 2008-A41
C/2007 T10  11.071 12 51.0 − 8 01 C3/2 HS 2008-A50
C/2007 T11  13.456 13 05.7 − 8 36 C2  HS 2008-A50

COMET 29P/SCHWASSMANN-WACHMANN This comet appears to be undergoing an outburst in brightness, as

indicated by the following visual total-magnitude estimates: 2007 Dec. 9.73 UT, 14.5 (S. Yoshida, Gunma, Japan, 0.40-m reflector); 31.10, 12.9 (M. Goiato, Aracatuba, Brazil, 0.22-m reflector); 2008 Jan. 3.87, 12.2 (J. J. Gonzalez, Asturias, Spain, 0.20-m refl.); 4.43, 12.8 (Yoshida); 11.89, 11.8 (M. L. Paradowski, Poleski National Park, Poland, 0.20-m refl.); 13.83, 11.4 (Gonzalez); 14.90, 11.3 (V. Nevski, Vitebsk, Belarus, 0.3-m refl.).

2008 January 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green